Fermi National Laboratory


Accelerator Update

Wednesday August 20
The day shift began with the Tevatron (TeV) recovering from a quench, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

At 9:07, the TeV suffered a sector B1 quench due to a separator problem.

Operations established a 12x0 study store at 12:16 PM.

Operations put the TeV into shot setup at 1:11 PM.

A Booster expert began work on a power supply (ORBMP) at 2:30 PM. This held off MiniBooNE beam. The Main Injector (MI) ramp was turned off during this repair period so the Recycler could perform a study.

Operations established store 2932 at 3:16 PM with and initial luminosity of 2.7E31.

The Booster expert completed his ORBMP work at 3:44 PM. A few minutes later, Operations had beam to MiniBooNE and Pbar stacking.

At 9:55 PM, operators noticed that the LCW flow at sectors A1 and F4 were a bit low. They contacted an expert who suggested rebooting a controller at CUB. This didn't help. An expert came in and discovered a bad power supply at 11:23 PM.

Thursday August 21
The midnight shift began with CUB experts replacing a controller power supply, with Operation monitoring store 2932, with Pbar stacking, and with MiniBooNE taking beam.

CUB experts had the system fully operational by 12:47 AM.

A Debuncher quadrupole power supply (D:Q731) tripped off at 4:20 AM due to a magnet high temperature, Twenty minutes later, operators were able to reset the supply. An expert said an access would be needed to investigate the magnet, but this would wait until some later time.

Pbar resumed stacking at 4:48 AM.

At 5:00 AM, two Linac RF stations (LRF3 & 4) began tripping on phase alarms. This was intermittently holding off beam.

At 5:13 AM, a Linac RF station (KRF5) tripped off due to a spark. It reset okay.

Operations terminated store 2932 at 7:16 AM. Fifteen minutes later, Operations put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 7:58 AM.

A Linac RF station (KRF5) tripped off four times in ten minutes starting at 8:40 AM. It reset okay each time.

KRF5 tripped again 9:04 AM.

Operations established store 2934 at 9:51 AM with and initial luminosity of 4.0E31. Pbar resumed stacking.

At 9:57 AM, a Debuncher RF station (DRF1) tripped off. An expert investigated and found two blown fuses. He replaced the fuses and had the station running by 10:15 AM.

Booster RF stations 5 & 6 tripped off at 1:28 PM due to a vacuum burst. The stations reset okay.

An expert completed a Booster study at 4:28 PM.

At 22:17 PM, Pbar lost its stack, multiple Booster RF station tripped, and much of Linac tripped off. One TeV RF station (TRF3) tripped and luminosity dropped a bit because of that, but the store remained in the machine; the cause of all this was traced to a ComEd power glitch.

Operations reestablished beam in Linac, Booster, and to MiniBooNE at 10:42 PM.

Pbar stacking resumed at 11:32 PM.

Friday August 22
The midnight shift began with Operations monitoring store 2934, with Pbar stacking, and with MiniBooNE taking beam.

At 12:43 AM, an Accumulator RF station (ARF1) began tripping off. After some odd problems, the station began operating normally at 1:05 PM. Pbar resumed stacking.

At 4:13 AM, Linac suffered more intertank phase problems with RF stations 3 & 4. Operators had the stations working by 4:30 AM.

At 6:53 AM, the Fermilab accelerators had delivered 225 pb-1 of integrated luminosity to the Collider Detector Facility (CDF) and D-zero (D0) experiments, which was our goal for this fiscal year.

Today's Plan
The plan for today is to terminate the store 2934 around 8:30 AM and then allow D0 and CDF to access their collision halls. After the experiments finish their access, TeV experts will conduct studies for six to eight hours. The weekend will consist of stacks and stores.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/22/2003   email Fermilab